A 121-kg astronaut (including space suit) acquires a speed of 2.20 m/s by pushing off with her legs from a 1850-kg space capsule. Use the reference frame in which the capsule is at rest before the push.

What is the velocity of the space capsule after the push in the reference frame?


If the push lasts 0.630 s , what is the magnitude of the average force exerted by each on the other?
Answer #1

a)

By Conservation of momentum

MAVA+MCVC=0

121*2.2+1850VC=0

VC=-266.2/1850 =-0.144 m/s

b)

Average force

Favg=mdV/dt =1850*0.144/0.63

Favg=422.54 N